A bond has a face value of $800 and a 10% coupon rate, its current price is S740, and it is expected to increase to $750 next year. The current yield is % (Enter your response rounded to one decimal place.) The expected rate of capital gain s %. Enter your response r unded to one decimal place) The expected rate of return is 1% (Enter your response rounded to one decimal place)
Answer #1

Answer : Current yield on bond = Annual interest payment/ Current market price

Annual interest payment = $800*0.10= $80

Market price = $740

Current yield on bond= $80/$740= 0.1081*100= 10.81%

Expected rate of capital gain =( ( New price - Old price) / Old price)*100= ((750-740)/740)*100= 1.35%

Expected rate of return on bonds = Current yield + Rate of capital gain = 10.81% + 1.35%= 12.16%